an electron (me = 9.10939 × 10–28 g) moving at 4.20 × 106 m/s in an electron microscope


1
Expert's answer
2022-03-27T10:27:33-0400

λ = h / p = h / m x v = 6.62x10-34 / 9.10939x10-28 x 3.9x106 = 1.8633914x10-13 m = 0.00018633914 nm